# Cosmology using numerical relativity

### v1

@@ 5 Revelation: The late universe
There exist several codes that are half-way houses between the Poisson-Vlasov approach and full NR, which are worth highlighting since they capture many of the potential relativistic effects with fewer overheads than a full NR simulation. For example, in gevolution [280], the Poisson equation (which arises from the $t-t$ component of the Einstein equations in the non-relativistic weak-field limit), is extended by the full $t-t$ and $i-j$ Einstein equations up to quadratic order, to better capture the dynamics of the gravitational degrees of freedom. The first equation is an elliptic equation that captures the dynamics of the Newtonian potential, while the second captures the tensor degrees of freedom as a hyperbolic equation. Another code is GRAMSES [281, 282], in which the metric degrees of freedom are obtained by solving the momentum and Hamiltonian constraints in-between time steps (instead of using the Poisson equation). This stratagem captures the gravitational degrees of freedom in the non-relativistic limit. In both codes, the matter sector is modelled by collisionless particles that move according to their corresponding geodesic equations.
@@ 5.1 Backreaction and averaging in non FLRW spacetimes
The first systematic investigations using NR to study late-time cosmological evolution were done using the CosmoGRaPH code of Giblin, Mertens & Starkman [208, 79], and in parallel by Bentivegna & Bruni [84] using the McLachlan evolution thorn of Einstein Toolkit suite of tools [283, 284]. See Fig. 19.
@@
In order to match numerical simulations to the FLRW quantities, the authors defined volume averaged quantities $\langle X\rangle_{S}$ over the spatial hyperslices $S$
@@
${\cal{S}}=\frac{2}{3}\left(\langle K^{2}\rangle-\langle K\rangle^{2}-2\langle A^{2}\rangle\right)~{},$ (70)
@@
measures the back-reaction from non-linear evolution. They found that for small values of initial perturbations, ${\cal{Q}}_{D}\propto a_{\cal{D}}^{-1}$, which agrees with 2nd order perturbation theory calculations [287]. They checked the validity of the standard semi-analytical model of collapse, the spherical top-hat model, in which a top-hat shaped overdensity is allowed to evolve under global expansion in a matter-dominated universe. The competition between infall and expansion leads to a moment of turn around when the gravitational collapse overcomes the local expansion. The critical density at turnaround $\delta_{OD}\equiv\delta\rho/\rho$ at the turn-around time $t_{TA}$. They found that $\delta_{OD}$ differed to the semi-analytic estimate of $\delta_{OD}\approx 1.06$, but a later work by Munoz & Bruni [288, 289] resolved this difference, finding $\delta_{OD}=1.055$ in good agreement with the semi-analytic results, thus validating the standard collapse model.
@@
The main results of these two works were subsequently confirmed by Macpherson et al. [290], who also used the Einstein Toolkit but incorporated the GRHydro thorn for more general matter and developed a new FLRWSolver thorn to generate the initial data and diagnostics. In this and later works [291, 292] Macpherson et al. simulated late-time structure formation in a matter-dominated universe, using a realistic matter power spectrum of perturbations derived from observations of the CMB as initial conditions. Due to limited resolution, they chose to simulate cosmological scales ranging from $k\sim 10^{-2}~{}\mathrm{Mpc}^{-1}$ to $k\sim 1~{}\mathrm{Mpc}^{-1}$. By implementing an averaging scheme which depends on the scale $r_{D}$, they found that while large-scale spatial homogeneity and isotropy is maintained, at scales $r_{D}\leq 180h^{-1}~{}\mathrm{Mpc}$, which correspond to local galaxy survey scales, deviation from homogeneity could be up to $6\%$. Nonetheless, this deviation only affects the measured value of the Hubble constant at the sub-percent level, and so is not sufficient on its own to resolve the current tension in its value, although in Macpherson [293] the value is found to be more significant for the case of observers located close to inhomogeneities. A study that added tensor perturbations was carried out by Wang [294].
@@
The codes used in these works undertook a comparison exercise [295] together with gevolution and GRAMSES. All codes ran a simulation with identical initial conditions and similar resolutions and compared the paths of two null, spatially separated trajectories, launched from the initial hyperslice. As the null trajectory is sensitive the entire spacetime metric, this serves as a good test of equivalence in relativistic cases. All codes showed excellent agreement (within $0.03\%$).
@@
A direct comparison between NR and N-body simulations was carried out by East et al. [296]. In this work, a full hydrodynamical treatment was used to model the collisionless dark matter [297], while the metric variables were evolved using the GHC formulation [90]. Meanwhile, GADGET-2 [3] was used to generate results for a non-relativistic N-body simulation for comparison to the NR results. Instead of computing deviations from linear theory using averaged quantities, the authors chose a set of fiducial observers that were co-moving with the matter, and tracked the evolution of $\delta\rho$ over their worldlines. They compared their results to both the Poisson-Vlasov results and expectations from linear perturbation theory. Furthermore, they solved for the geodesics of null rays, and therefore allowing them to compute the cosmological distance as a function of redshift for free-falling observers located along the photon paths. They found that, as long as the Newtonian potential remains in the weak-field limit, the difference between the NR simulations to both the N-body simulations and linear theory expectation is at the sub-percent level, see Fig. 20.
@@
The use of NR to probe the dynamics of expanding universes and the impact of backreaction in inhomogeneous cases has also been tackled from the perspective of a lattice of BHs. Roughly speaking, if the scale of the BHs is much shorter than the curvature scale of the universe $M_{BH}\ll H^{-1}$, then a lattice of static BHs mimics (at the largest scales) that of cold dark matter (see e.g. [302]), which can drive cosmic expansion. Thus such a spacetime is an interesting toy model, with which one can study the physics of dark matter and the role of averaging in FLRW, at least on the largest scales. This was first proposed by Linquist and Wheeler [303], and then numerically simulated in several works [304, 305, 306, 307, 308, 309, 310], see Bentivegna et al. [311] for a review. These simulations have studied effects such as the dressing (differences from the naive superposition of constituents) of mass and other physical parameters, and the biases in optical measurements that inhomogeneous gravitational fields can produce.
@@ 5.2 Cosmological observables in the non-linear regime
A general theme of the results described in the previous works in this section is that the Poisson-Vlasov formalism is remarkably accurate in capturing most of the relevant gravitational effects in late-time cosmological evolution, even in the non-linear regime. Significant deviations from full NR do not occur until deep in the strong gravity regime – expected at very short wavelength scales when local regions decouple fully from the Hubble flow. Indeed, these results indicate that scale-separation is robust in cosmological evolution – non-linearity of GR does not induce UV-IR couplings where UV modes significant affect large-scale cosmological evolution.
@@
Nevertheless, as mentioned earlier, NR still has a role to play both in terms of understanding the systematics of $N$-body predictions and as a useful cross-check. The first such cross-check was undertaken by Giblin et al. [313] where they computed the weak lensing power spectrum using NR. Weak lensing power computations often rely on approximations [314]. In particular, line-of-sight integrals often rely on the so-called Born approximation, where the light ray is assumed to propagate along the background geodesic and is not perturbed by higher order gravitational effects. Attempts to incorporate post-Born effects has led to disagreement [^13] Although [315, 316] argued that the disagreement is not related to the Born approximation (which still applies) but a disagreement in the labeling of what constitutes lens-lens couplings. [317, 318]. Giblin et al. [313] attempted to resolve this issue by computing the weak lensing power spectrum by ray-tracing a universe evolved using NR. They found that the approximation may lead to percent level error at angular scales of $l=20\sim 30$ when comparing full NR to perturbation theory computations, potentially signalling important systematics not captured in higher order perturbation theory. More recent efforts to develop tools to probe cosmological distances using ray tracing have been carried out [319, 320], and support the conclusion that effects may be at percent level for scales above 100 Mpc.
@@
Macpherson & Heinesen [321] evolved an anisotropic cosmological spacetime, and then computed observables such as the Hubble parameter and its higher moments with different levels of sky coverage. They showed that there was up to percent level cosmic variance in the measurements. For example, there is a $7\%$ probability that an observer can measure up to a $3\%$ difference in the Hubble parameter of a typical Einstein-de-Sitter universe, see Fig. 21. In related works, Macpherson [293] and Williams et al. [322] have studied the impact of nearby inhomogeneities and voids on observers’ inferences of the Hubble constant, where they found that the voids are in general negatively curved and void statistics such as size distribution roughly follows that of $N$-body simulations. Koksbang et al. [323] extended this by computing the red-shift drift – where the redshift of an object evolves through time due to cosmic expansion. They showed that NR predicts that at redshift $z=0.1$, fluctuations in the redshift drift due to inhomogeneities are at the $\sim 10\%$ level, and so potentially observable in the next generation spectroscopic sky surveys.
@@ 7 Summary table and open questions
Bounces [160, 169, 161, 163, 165, 168] [165] (fully 3D),[164, 53] (axisymmetric) [160, 161, 163, 164, 165, 168, 169] [53] – – [160, 161, 163, 164, 165, 168, 169, 53] – – [169, 53] [160, 161, 163, 164, 165, 168]
@@
    How does gravity and AMR impact the evolution of a network of early universe relics? e.g. oscillons, topological defects.
@@
Revelation: The late universe
@@
    What is the impact of including GR effects on systematics for next generation cosmological surveys?
@@ Acknowledgements
We acknowledge helpful input from many of the authors of the papers presented here. JCA acknowledges funding from the Beecroft Trust and The Queen’s College via an extraordinary Junior Research Fellowship (eJRF). KC acknowledges funding from from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 693024), an STFC Ernest Rutherford Fellowship project reference ST/V003240/1 and STFC Research Grant ST/X000931/1 (Astronomy at Queen Mary 2023-2026). EAL acknowledges support from a Leverhulme Trust Research Project Grant.

### v2

@@ 5 The late universe
There exist several codes that are half-way houses between the Poisson-Vlasov approach and full NR, which are worth highlighting since they capture many of the potential relativistic effects with fewer overheads than a full NR simulation. For example, in gevolution [21], the Poisson equation (which arises from the $t-t$ component of the Einstein equations in the non-relativistic weak-field limit), is extended by the full $t-t$ and $i-j$ Einstein equations up to quadratic order, to better capture the dynamics of the gravitational degrees of freedom. The first equation is an elliptic equation that captures the dynamics of the Newtonian potential, while the second captures the tensor degrees of freedom as a hyperbolic equation. Another code is GRAMSES [23, 24], in which the metric degrees of freedom are obtained by solving the momentum and Hamiltonian constraints in-between time steps (instead of using the Poisson equation). This stratagem captures the gravitational degrees of freedom in the non-relativistic limit. In both codes, the matter sector is modelled by collisionless particles that move according to their corresponding geodesic equations. Much of the NR work described below has used pressureless (or very low pressure) fluids to model the dark matter evolution, building on existing NR capabilities. This is not ideal since the description and imposition of a fixed equation of state breaks down once collapsed structures begin to form. For precision work, hybrid NR-N-body codes that account for kinetic pressure and permit virialisation are required. However, such studies have provided a useful first step in looking at the impact of general relativistic effects on large cosmological scales.
@@ 5.1 Backreaction and averaging in non FLRW spacetimes
The first systematic investigations using NR to study late-time cosmological evolution were done using the CosmoGRaPH code of Giblin, Mertens & Starkman [257, 118], and in parallel by Bentivegna & Bruni [123] using the Einstein Toolkit [334, 335], closely (and independently) followed by Macpherson, Lasky & Price [336] also using the Einstein Toolkit.
@@
In order to match numerical simulations to the FLRW quantities, the authors defined volume averaged quantities $\langle X\rangle_{S}$ over the spatial hyperslices $S$ [328]
@@
${\cal{Q}_{\cal{S}}}=\frac{2}{3}\left(\langle\Theta^{2}\rangle-\langle\Theta\rangle^{2}-2\langle\sigma^{2}\rangle\right)~{},$ (72)
@@
measures the back-reaction from non-linear evolution, with $\Theta$ the expansion of the fluid and $\sigma^{2}=\sigma_{ij}\sigma^{ij}$ its shear squared. They found that for small values of initial perturbations, ${\cal{Q}}_{D}\propto a_{\cal{D}}^{-1}$, which agrees with 2nd order perturbation theory calculations [339]. Using a simplified model (a sine wave), they checked the validity of the standard model of collapse, the spherical top-hat model, in which a top-hat shaped overdensity is allowed to evolve under global expansion in a matter-dominated universe. The competition between infall and expansion leads to a moment of turn around when the gravitational collapse overcomes the local expansion. The critical density at the turn-around time $t_{TA}$ is $\delta_{OD}\equiv\delta\rho/\rho$. They found that $\delta_{OD}$ differed from the estimate of $\delta_{OD}\approx 1.06$, due to the presence of a growing and decaying mode in the initial data (because of the zero velocity assumption, as pointed out by East et al. [340]). A later work by Munoz & Bruni [148] resolved this difference by fixing the initial data, finding $\delta_{OD}=1.055$ in good agreement with the semi-analytic results, thus validating the standard collapse model.
@@
The main results of these two works were subsequently confirmed by Macpherson et al. [336, 341], who also used the Einstein Toolkit but incorporated the GRHydro thorn for more general matter (allowing the addition of a small pressure to the dust model) and developed a new FLRWSolver thorn to generate the initial data, along with separate diagnostics tools. In this and later works [341, 342] Macpherson et al. simulated late-time structure formation in a matter-dominated universe, using a realistic matter power spectrum of perturbations derived from observations of the CMB as initial conditions. Due to limited resolution, they chose to simulate cosmological scales ranging from $k\sim 10^{-2}~{}\mathrm{Mpc}^{-1}$ to $k\sim 1~{}\mathrm{Mpc}^{-1}$. By implementing an averaging scheme which depends on the scale $r_{D}$, they found that while large-scale spatial homogeneity and isotropy is maintained, at scales $r_{D}\leq 180h^{-1}~{}\mathrm{Mpc}$, which correspond to local galaxy survey scales, deviation from homogeneity could be up to $6\%$. Nonetheless, this deviation only affects the measured value of the Hubble constant at the sub-percent level, and so is not sufficient on its own to resolve the current tension in its value, although in Macpherson [343] the value is found to be more significant using ray tracing instead of spatial averages. A study that added tensor perturbations was carried out by Wang [344].
@@
The codes used in these works undertook a comparison exercise [22] together with gevolution and GRAMSES. All codes ran a simulation with identical initial conditions (a plane wave) and similar resolutions and compared the paths of two null, spatially separated trajectories, launched from the initial hyperslice. As the null trajectory is sensitive the entire spacetime metric, this serves as a good test of equivalence in relativistic cases. All codes showed excellent agreement (within $0.03\%$).
@@
A direct comparison between NR and N-body simulations was carried out by East et al. [340]. In this work, a full hydrodynamical treatment was used to model the collisionless dark matter [345], while the metric variables were evolved using the GHC formulation [130]. Meanwhile, GADGET-2 [5] was used to generate results for a non-relativistic N-body simulation for comparison to the NR results. Instead of computing deviations from linear theory using averaged quantities, the authors chose a set of fiducial observers that were co-moving with the matter, and tracked the evolution of $\delta\rho$ over their worldlines. They compared their results to both the Poisson-Vlasov results and expectations from linear perturbation theory. Furthermore, they solved for the geodesics of null rays, and therefore allowing them to compute the cosmological distance as a function of redshift for free-falling observers located along the photon paths. They found that, as long as the Newtonian potential remains in the weak-field limit, the difference between the NR simulations to both the N-body simulations and linear theory expectation is at the sub-percent level, see Fig. 20. Their results were also found to be consistent with the spherical top-hat model of collapse.
@@
The use of NR to probe the dynamics of expanding universes and the impact of backreaction in inhomogeneous cases has also been tackled from the perspective of a lattice of BHs. Roughly speaking, if the scale of the BHs is much shorter than the curvature scale of the universe $M_{BH}\ll H^{-1}$, then a lattice of static BHs mimics (at the largest scales) that of cold dark matter (see e.g. [350]), which can drive cosmic expansion. Thus such a spacetime is an interesting toy model, with which one can study the physics of dark matter and the role of averaging in FLRW, at least on the largest scales. This was first proposed by Linquist and Wheeler [351], and then numerically simulated in several works [352, 353, 354, 355, 356, 357, 358], see Bentivegna et al. [359] for a review. These simulations have studied effects such as the dressing (differences from the naive superposition of constituents) of mass and other physical parameters, and the biases in optical measurements that inhomogeneous gravitational fields can produce. They find that as the number of BHs increase, backreaction decreases.
@@ 5.2 Cosmological observables in the non-linear regime
The results of [346] described in the previous section showed that the Poisson-Vlasov formalism is remarkably accurate in capturing most of the relevant gravitational effects in late-time cosmological evolution, even in the non-linear regime. Significant deviations from this approximate formalism as compared to a fully relativistic treatment do not occur until deep in the strong gravity regime – expected at very short wavelength scales when local regions decouple fully from the Hubble flow. Indeed, these results indicate that scale-separation is robust in cosmological evolution – non-linearity of GR does not induce UV-IR couplings where UV modes significantly affect large-scale cosmological evolution, at least not sufficiently to explain the observed acceleration of the expansion of the universe. However, they cannot entirely rule out percent level effects that may be be relevant for future precision cosmological observations, for example galaxy surveys.
@@
NR still has a role to play both in terms of understanding the systematics of $N$-body predictions and as a useful cross-check. The first such cross-check was undertaken by Giblin et al. [361] where they computed the weak lensing power spectrum using NR. Weak lensing power computations often rely on approximations [362]. In particular, line-of-sight integrals often rely on the so-called Born approximation, where the light ray is assumed to propagate along the background geodesic and is not perturbed by higher order gravitational effects. Attempts to incorporate post-Born effects has led to disagreement [^18] Although [363, 364] argued that the disagreement is not related to the Born approximation (which still applies) but a disagreement in the labeling of what constitutes lens-lens couplings. [365, 366]. Giblin et al. [361] attempted to resolve this issue by computing the weak lensing power spectrum by ray-tracing a universe evolved using NR. They found that the approximation may lead to percent level error at angular scales of $l=20\sim 30$ when comparing full NR to perturbation theory computations, potentially signalling important systematics not captured in higher order perturbation theory. More recent efforts to develop tools to probe cosmological distances using ray tracing have been carried out [367, 368], and support the conclusion that effects may be at percent level for scales above 100 Mpc.
@@
Macpherson & Heinesen [369] studied the anisotropic effects due to small-scale structure, from initial data that were statistically homogeneous and isotropic, and computed observables such as the Hubble parameter and its higher moments with different levels of sky coverage. They found up to percent level cosmic variance in the measurements. For example, there is a $7\%$ probability that an observer can measure up to a $3\%$ difference in the Hubble parameter of a typical Einstein-de-Sitter universe, see Fig. 21. In related works, Macpherson [343] and Williams et al. [370] have studied the impact of nearby inhomogeneities and voids on observers’ inferences of the Hubble constant, where they found that the voids are in general negatively curved and void statistics such as size distribution roughly follows that of $N$-body simulations. Koksbang et al. [371] extended this by computing the red-shift drift – where the redshift of an object evolves through time due to cosmic expansion. They showed that NR predicts that at redshift $z=0.1$, fluctuations in the redshift drift due to inhomogeneities are at the $\sim 10\%$ level, and so potentially observable in the next generation spectroscopic sky surveys. Another observable of interest is frame dragging, arising from cosmological vector modes caused in the late universe by non-linearities. Frame dragging is the leading-order correction in a post-Newtonian formalism and is expected to be generated after the first shell-crossing in nonlinear structure formation, even in the weak-field regime [372, 373, 374, 375]. Some initial work using the relativistic codes GRAMSES and gevolution has been done in this direction [21, 376, 377] and the first steps using NR are taken in [143, 148].
@@ 7 Summary table and open questions
Slow contraction/Bounces [209, 218, 210, 212, 214, 217] [214] (fully 3D),[213, 85] (axisymmetric) [209, 210, 212, 213, 214, 217, 218] [85] – – [209, 210, 212, 213, 214, 217, 218, 85] – – [218, 85] [209, 210, 212, 213, 214, 217]
@@
    How does gravity and accurately resolving the small-scale dynamics of early universe relics (e.g. the cosmic string core) affect the evolution of the network?
@@
  6. 6.
    How does the GW spectrum of a cosmological phase transition change when studying the dynamics of the gauge theory using holography?
The late universe
@@
    What is the impact of including GR effects on systematics and percent level effects that may arise in next generation cosmological surveys? e.g. frame dragging, spatial curvature, voids and overdensities, red-shift drift.
@@
    How does one disentangle which effects are “backreaction” effects, and which are numerical systematics such as differences in methodologies?
  3. 3.
@@ Acknowledgements
We acknowledge helpful input from many of the authors of the papers presented here, in particular Thomas Baumgarte, Marco Bruni, Maxence Corman, Will East, David Garfinkle, Tom Giblin, Sofie Marie Koksbang, Hayley Macpherson and Robyn Munoz. We also thank the anonymous referees for their comments, which helped us improve this manuscript. JCA acknowledges funding from the Beecroft Trust and The Queen’s College via an extraordinary Junior Research Fellowship (eJRF). KC acknowledges funding from from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 693024), an STFC Ernest Rutherford Fellowship project reference ST/V003240/1 and STFC Research Grant ST/X000931/1 (Astronomy at Queen Mary 2023-2026). EAL acknowledges support from a Leverhulme Trust Research Project Grant.